Grade 12General Physics

The retardation experienced by a moving motor boat after its engine is cut off is given by:dv/dt=-kv3 ,where k is a +ve constant.If v0 is the magnitude of velocity at cut-off,the magnitude of the velocity at time t after the cut-off is:

Vaishal Shah
15 Years agoGrade 12

2 Answers

Chetan Mandayam Nayakar
15 Years ago

dv/v^3=-kdt,(v^-2)/(-2)=-kt+c,c=-(1/2v2)

KARAN ACHARYA
15 Years ago

dv/dt = -kv.v.v

dv/(kv.v.v) = -dt

integrate[v Limit v0 to v    t Limit 0 to t]

1/[v.v] = 1/[v0.v0] +2kt

v^2  = [(1/v0.v0) +2kt]^[-2]

v = [(1/v0.v0) +2kt]^[-1]